# Joint Liability Model with Adaptation to Climate Change

### v1

@@ 5 Conclusions
In this paper, we incorporate the individual-ESE score into a joint liability model for evaluating the sustainability and creditworthiness of individual farmers acting as borrowers in the model. To achieve this objective, we initially only considered the expected profit of an individual borrower resulting from the specification of his/her utility function. Our primary objective was to determine the optimal group size that maximizes the benefit for borrowers while ensuring that the lender breaks even. Additionally, we sought to derive the optimal Individual-ESE Score, which can provide several benefits to both lending companies and farmers as borrowers, such as a better understanding of the borrower’s risk profile, information about the optimal group size for lending, and improved access to credit. The inclusion of this score in credit evaluation processes can lead to better-informed lending decisions and sustainable development outcomes. We listed five propositions. In this paper, proposition 1 focused on finding the optimal individual-ESE score. Proposition 2 aimed at reducing the two constraints about the loan ceiling to only one constraint. Proposition 3 demonstrated that a higher individual-ESE score leads to a higher loan ceiling. Proposition 4 revealed that the optimal individual-ESE score and group size n have a negative relationship. Finally, Proposition 5 showed that the optimal individual-ESE score converges to a constant level when the group size reaches a large number.
Furthermore, our study has also incorporated the risk associated with expected profit into the utility function of the farmers and analyzed the resulting mean-variance utility framework in the context of joint liability problems. The study additionally disentangled the probability of achieving success in a project into two constituent components, namely the impact of personal factors (represented by $k\times ESE$) and exogenous factors beyond individual farmer’s control (represented by $b$), such as climatic and geographical conditions. Our findings reveal that the probability of a project’s success is inversely proportional to the required ESE score. This implies that a higher likelihood of project success leads to increased farmers’ ability to repay their loans, thereby lowering the risk of default for the insurance company. As a result, the insurance company can lower its ESE score requirements when lending to high-quality farmers. We have also found that an increase in farmers’ risk aversion score leads to more cautious contract behavior, reducing the probability of strategic default and increasing the probability of repayment. Hence, as the risk aversion score increases, the corresponding ESE score requirement decreases. Finally, an increase in the cost parameter raises the production expenditure of farmers and burdens their economic resources, reducing their profits. Therefore, to avoid the high default probability associated with high costs, the corresponding ESE score requirement should increase.
Subsequently, we proceed to explore the effects of climate change on the required ESE score in order to ensure a successful project implementation. Our simulation results demonstrate that a higher likelihood of project success, driven by high yields resulting from favorable climatic conditions, leads to a lower ESE score requirement in a joint liability contract. This suggests that incorporating climate scenarios into the evaluation of creditworthiness can result in better-informed lending decisions and ultimately contribute to more favorable sustainable development outcomes.
As an avenue for future research, this study paves the way for exploring partial joint liability – a legal framework distributing shared responsibility among parties for specific obligations. Unlike complete joint liability, partial joint liability assigns a proportionate debt share to each party. Moreover, the concept of flexible joint liability introduces innovation. Participants begin with collaborative lending similar to standard joint liability but, in strategic default scenarios, opt for a punitive phase rather than immediate maximum penalties. After a defined period ($T$ as previously mentioned), defaulters are reintegrated, offering a path for forgiveness and adherence to the flexible joint liability contract. The diverse joint liability arrangements highlight the need for a thorough and separate analysis. Last but not least, our review of the literature has revealed that while traditional credit metrics concentrate on financial and economic factors, assessing sustainability requires taking into account environmental and social factors as well. Therefore, we have suggested the concept of an individual-ESE score that integrates economic, social, and environmental considerations to offer a more all-encompassing appraisal of the farming sector. It follows that another avenue for future research is to eventually develop a credit scoring system that addresses the unique requirements of the agricultural sector by applying literature-based metrics for agricultural credit scoring systems to the construction of the scores. By compiling and adjusting these metrics, we eventually aim at creating a comprehensive credit scoring system with adaptation to climate change for the agricultural sector that addresses limitations of the ESG scores when they are applied to the agricultural sector.

### v2

@@ 4 Conclusions
In this paper we have incorporated a novel concept coined as an individual–ESE score into a joint liability model for the purpose of evaluating the sustainability and creditworthiness of individual farmers acting as borrowers in the model. To achieve this objective we initially considered the expected profit of an individual borrower resulting from the specification of his/her utility function. Our primary objective was to determine the optimal group size that maximizes the benefit to borrowers while ensuring that the lender breaks even. Additionally we sought to derive the optimal individual–ESE score, which can provide several benefits to both the lending companies and the farmers as borrowers, such as a better understanding of the borrower’s risk profile, information about the optimal group size for lending, and an improved access to credits. The inclusion of this score in the credit evaluation processes can lead to better–informed lending decisions and sustainable development outcomes. We listed the main results of this paper conveniently in five propositions. Proposition 1 focused on finding an optimal individual-ESE score. Proposition 2 aimed at collapsing two constraints on the loan ceiling to a single constraint. Proposition 3 demonstrated that a higher individual-ESE score leads to a higher loan ceiling. Proposition 4 revealed that the optimal individual-ESE score and group size $n$ have a negative relationship. Finally Proposition 5 showed that the optimal individual–ESE score converges to a constant level when the group size reaches a large enough number. As a next step our study incorporated the risk associated with an expected profit into the utility function of the farmers and analyzed the resulting mean-variance utility framework in the context of the individual $ESE$–joint liability model. The study additionally disentangled the probability of achieving success in a project into two constituent components, namely the impact of personal factors (represented by $k\times ESE$) and exogenous factors beyond individual farmer’s control (represented by $b$), such as climatic and geographical conditions. Our findings reveal that the probability of a project’s success is inversely proportional to the required ESE score. This implies that a higher likelihood of a project success leads to an increased farmers’ ability to repay their loans, thereby lowering the risk of default for the insurance company. As a result the insurance company can lower its $ESE$ score requirements when lending to high-quality farmers. We also found that an increase in farmers’ risk aversion score leads to a more cautious contract behavior, reducing the probability of a strategic default and increasing the probability of repayment. Hence, as the risk–aversion score increases, the corresponding $ESE$ score requirement decreases. Finally an increase in the cost parameter raises the production expenditure of the farmers and places a burden on their economic resources which reduces their profits. Therefore, to avoid a high default probability associated with high costs, the corresponding requirement of an ESE score should be increased. Subsequently we explored the effects of climate change on the required ESE score in order to ensure a successful project implementation. Our simulation results demonstrated that a higher likelihood of a project success, driven by high yields resulting from favorable climate conditions, leads to a lower $ESE$ score requirement in the proposed individual ESE–joint liability model. This suggests that incorporating climate scenarios into the evaluation of creditworthiness can result in better–informed lending decisions and ultimately contribute to more favorable sustainable development outcomes. As an avenue for future research, this study paves the way for an exploration of a partially joint liability’s legal framework which distributes shared responsibility among parties for specific obligations. Unlike a complete joint liability frameworrk, a partially joint liability framework assigns a proportionate debt share to each party. Moreover the concept of a flexible joint liability introduces aspects of innovation. Participants begin with a collaborative lending similar to a standard joint liability contract but, in strategic default scenarios, opt for a punitive phase rather than immediate maximum penalties. After a defined period ($T$ as previously mentioned), defaulters are reintegrated, which offers a path for forgiveness and adherence to the flexible joint liability contract. The diverse joint liability arrangements highlight the need for a thorough and separate analysis. Last but not least, our review of the literature has revealed that while traditional credit metrics concentrate on financial and economic factors, assessing sustainability requires taking into account environmental and social factors as well. Therefore, as a first step of analysis, we have suggested the concept of an individual-ESE score that integrates ESE considerations to offer a more encompassing appraisal of the farming sector. It follows that a potentially most consequential avenue for a future research is to eventually develop a credit scoring system that addresses the unique requirements of the agricultural sector by applying literature–based metrics for agricultural credit scoring systems to the construction of the scores. By compiling and adjusting these metrics, we hope to eventually be able to create a comprehensive credit scoring system with adaptation to climate change for the agricultural sector that addresses limitations of the ESG scores as they are applied to the agricultural sector. This is an important task to take up but is clearly beyond the scope of the present study.